OpenStudy (anonymous):

What is the measure of one interior angle in a regular hexagon? 135° 120° 360° 720°

OpenStudy (anonymous):

\[\frac{ (n - 2) * 180° }{ n }\] hexagon has 6 sides, so n = 6

OpenStudy (tkhunny):

EXterior angles are ALWAYS 360º. Hexagon? 6 of them. 360º / 6 = 60º <== Exterior Interior and Exterior angles form Linear Pairs. 180º - 60º = 120º <== Interior!